OpenStudy (anonymous):

Solve the equation for x. –x + 10 – 8x = 17 – 2x A.11 B.–11 C.1 D.–1

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

show me your steps

OpenStudy (anonymous):

someone else did it for me idk how they did it tho?

OpenStudy (anonymous):

ok I will show you

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

–x + 10 – 8x = 17 – 2x <----- Combine like terms -8x -x = -9x -9x + 10 = 17 - 2x <-- now move -2x to the left by adding it, add it to -9x +2x +2x ------------------ -7x + 10 = 17 <----- now move +10 to right by subtracting it, also subtract to 17 -10 -10 -7x = 7 < ----- now divide both sides by -7 ------- -7 -7 x = -1

OpenStudy (anonymous):

that's not how the other person showed me, thanks that makes way more sense.

OpenStudy (anonymous):

theres different ways, but this is the easiest I think

OpenStudy (anonymous):

do you want me to give you an easy problem? so you can try it?